# Symmetry-Aware GFlowNets

### v1

@@ 3.2 Generative Flow Networks
$\vspace{0pt}{\mathcal{L}}_{\mathrm{TB}}(\tau)=\left(\log\frac{Z\prod_{t=0}^{n-1}p_{\mathcal{A}}(s_{t+1}|s_{t})}{R(s_{n})\prod_{t=0}^{n-1}q_{\mathcal{A}}(s_{t}|s_{t+1})}\right)^{2}.$
@@ 4.1 Problem Definition
Since isomorphism is an equivalence relation, it partitions the space ${\mathcal{G}}$ into classes, where each graph in a class is structurally identical to the others. Let $[G]=\{G^{\prime}\in{\mathcal{G}}:G^{\prime}\cong G\}$ denote the equivalence class of $G$ induced by graph isomorphism. The state space ${\mathcal{S}}$ is defined as the set of equivalence classes of graphs, ${\mathcal{S}}=\{[G]:G\in{\mathcal{G}}\}$, rather than the graph space ${\mathcal{G}}$ itself. This is because our goal in using GFlowNets is to sample any graph within the equivalence class $s=[G]$ in proportion to $R(s)$. State transitions ${\mathcal{A}}$ can also be defined by partitioning the graph transitions ${\mathcal{E}}$ by the following equivalence relation:
@@
In practice, a graph generation process $({\mathcal{G}},{\mathcal{E}})$ is constructed by first designing a set of allowable actions in a given graph. For example, AddEdge$(G,u,v)$ adds an edge $(u,v)$ to the existing graph $G$, and $\texttt{AddNode}(G,u)$ adds a new node to an existing node $u$. The $\texttt{Stop}(G)$ terminates the process, in which case the graph-level attribute is flagged as terminated. Through this construction, graph transitions ${\mathcal{E}}$ are ensured to be structured so that any pair of isomorphic graphs have isomorphic successors.
@@
As an alternative, Ma et al. (2024) suggested using orbits to identify transition-equivalent actions, which we formalize next. Define graph actions as triples, ¯=$\bar{\mathcal{E}}=\{(G,t,u)\}$, where $t$ denotes the action type and $u$ specifies the nodes or edges affected by the action. Since graph transitions ${\mathcal{E}}$ are constructed by predefined set of graph actions, there is a one-to-one correspondence between ${\mathcal{E}}$ and $\bar{\mathcal{E}}$. However, we explicitly distinguish between them to introduce another equivalence relation.
@@
Given the distinction between transition-equivalent actions ${\mathcal{A}}$ and orbit-equivalent actions $\bar{\mathcal{A}}$, we define the state-action probability $p_{\bar{\mathcal{A}}}(a|s)$ differently from the state transition probability $p_{\mathcal{A}}(s^{\prime}|s)$: while $p_{\mathcal{A}}(s^{\prime}|s)$ accounts for all transition-equivalent actions, $p_{\bar{\mathcal{A}}}(a|s)$ aggregates only over orbit-equivalent actions. Formally, for a given graph $G$, let ¯$\bar{\mathcal{E}}(G)$ denote the set of graph actions available from $G$. Then, for a state-action pair $s\in{\mathcal{S}}$ and $a\in\bar{\mathcal{A}}$, the state-action probability is defined as:
@@
where $p_{\bar{\mathcal{E}}}(e|G)$ denotes $p_{\mathcal{E}}(G^{\prime}|G)$ for $G^{\prime}$ being the next graph. Here, ¯$\bar{\mathcal{E}}(G)\cap a$ represents the set of orbit-equivalent actions from $G$. By Theorem 4.3, it follows that $p_{\bar{\mathcal{A}}}(a|s)\leq p_{\mathcal{A}}(s^{\prime}|s)$ in general.
@@ 4.2 Properties of Equivalent Actions
Figure 2: Graphs representing transitions $(G_{1},G_{2},G_{3})$, where the first transition is performed by AddNode and the second by AddEdge. The number of foward/backward actions are represented as the number of arrows. Symmetries in each graph is related to orbit-equivalent actions, as seen in the ratio $|{\mathrm{Aut}}(G_{1})|/|{\mathrm{Aut}}(G_{2})|=|{\mathrm{Orb}}(G_{1},$ $)|/|{\mathrm{Orb}}(G_{2},$ )|)|) |. Nodes in the same orbit are given the same color.
@@
where |¯$|\bar{\mathcal{E}}(G)\cap a|$ represents the number of orbit-equivalent actions. This is because when actions are parameterized using graph neural networks (GNNs), orbit-equivalent actions are assigned equal probabilities. In general, permutation-equivariant functions produce identical representations for nodes within the same orbit (as detailed in Appendix E.1). When node representations are aggregated to compute edge representations using invariant aggregators such as SUM or MEAN, edges within the same orbit also receive identical representations. Alternative parameterizations, such as the relative edge parameterization proposed by Shen et al. (2023), also assign equal probabilities to orbit-equivalent actions, while enhancing representational power.
Note that the number of orbit-equivalent actions, |¯$|\bar{\mathcal{E}}(G)\cap a|$, corresponds to the size of the orbit of the associated nodes or edges. The following lemma shows that, when considering ratios, counting orbit-equivalent actions simplifies to counting automorphisms.
@@
We presented Lemma 4.5 in the context of AddEdge for simplicity, but similar lemma holds for other action types such as AddNode, AddNodeAttributes, and others used in node-by-node generation (See Appendix C for the full list of considered actions). We also consider AddFragment in our experiments, but discuss its properties separately in Appendix H.
@@ 5 Symmetry-Aware GFlowNets
Assume that $G_{0}$ is the empty graph or a single node, so that $|{\mathrm{Aut}}(G_{0})|=1$. Given the complete graph trajectory $\tau=(G_{0},G_{1},\dots,G_{n})$, the trajectory balance loss can be written as follows:
@@
We define the graph-level detailed balance condition, as opposed to the usual state-level condition, as follows:
@@
Together with Corollary 5.1, we see that scaling the reward alone is sufficient for both TB and DB objectives. This suggests that other GFlowNet objectives, such as subtrajectory balance (Madan et al., 2023) and flow-matching (Bengio et al., 2021), can also be used with reward scaling (see Appendix G). This provides a straightforward approach to implementing GFlowNet objectives while reducing the computational burden of counting automorphisms at each transition.
@@
Let $G$ represents a terminal state ($[G]\in{\mathcal{X}}$) generated by connecting $k$ fragments $\{C_{1},\dots,C_{k}\}$. Then, the scaled rewards to offset the effects of equivalent actions are given by:
@@ 6 Experiments
In this section, we conduct experiments to validate our theoretical results and demonstrate the effectiveness of our method. We use a uniform backward policy across all experiments. Details on hyperparameters and model configurations can be found in Appendix K. The experiments compare the following methods: 1) Vanilla GFlowNets, which do not incorporate graph symmetries. 2) Transition Correction, which identifies transition-equivalent actions by performing multiple isomorphism tests and sums their probabilities accordingly. 3) PE, the method proposed by Ma et al. (2024), which approximately identifies orbit-equivalent actions using positional encoding. 4) Reward Scaling, which achieves correction by modifying only the reward signals. 5) Flow Scaling, which multiplies symmetry ratio $|{\mathrm{Aut}}(G^{\prime})|/|{\mathrm{Aut}}(G)|$ to backward probability at each transition. Note that methods 3-5 are orbit correction methods. Since Reward Scaling and Flow Scaling have the same effect under the TB objective, we only consider Flow Scaling when using the DB objective. [^1] Source code available at: https://github.com/hohyun312/sagfn
@@ 6.3 Molecule Generation
Atom Vanilla 0.929$\pm$0.024 0.077$\pm$0.022 1.09$\pm$0.02 1.09$\pm$0.02 0.93$\pm$0.077
Reward Scaling (Exact) 0.959$\pm$0.01 0.046$\pm$0.006 1.091$\pm$0.013 1.091$\pm$0.013 1.0$\pm$0.0
Fragment Vanilla 0.877$\pm$0.001 0.153$\pm$0.003 0.941$\pm$0.002 0.941$\pm$0.002 1.0$\pm$0.0
Reward Scaling (Approx.) 0.88$\pm$0.001 0.164$\pm$0.008 0.949$\pm$0.006 0.949$\pm$0.006 1.0$\pm$0.0
Reward Scaling (Exact) 0.879$\pm$0.0 0.151$\pm$0.002 0.952$\pm$0.003 0.952$\pm$0.003 1.0$\pm$0.0
@@ 7 Discussion and Conclusion
In this paper, we analyzed the properties of equivalent actions and proposed a simple correction method that allows for unbiased sampling from the target distribution. Our analysis shows that, without correction, highly symmetric graphs are less likely to be sampled, while symmetric fragments are more likely to be sampled, which is crucial for molecule discovery. We demonstrated that the reward-scaling technique works for both TB and DB objectives. Experimental results suggest that reward scaling effectively removes bias, allowing for accurate modeling of the target distribution, which is essential for sampling high-reward molecules.
@@ Appendix C Definitions of Graph Actions
  • AddNode$(G,u)$ adds a new node to the existing node $u$.
  • AddEdge$(G,u,v)$ adds a new edge $(u,v)$.
  • AddFragment$(G,C)$ adds a fragment $C$.
  • RemoveNode$(G,v)$ removes node $v$ and its connecting edges.
  • RemoveEdge$(G,u,v)$ removes the edge $(u,v)$.
  • RemoveFragment$(G,C)$ removes the subgraph $C$.
  • SetNodeAttribute$(G,u,t)$ sets the node-level attribute $t$ for node $u$.
  • SetEdgeAttribute$(G,u,v,t)$ sets the edge-level attribute $t$ for the edge $(u,v)$.
  • SetGraphAttribute$(G,t)$ sets a graph-level attribute $t$.
@@ F.1 Proof of Equation 5
$\pi(E\cup E^{\prime})$ =$=\{(\pi(u),\pi(v)):(u,v)\in E\cup E^{\prime}\}$
=$=\{(\pi(u),\pi(v)):(u,v)\in E\text{ or }(u,v)\in E^{\prime}\}$
=$=\{(\pi(u),\pi(v)):(u,v)\in E\}\cup\{(\pi(u),\pi(v)):(u,v)\in E^{\prime}\}$
@@
$\pi(E\setminus E^{\prime})$ =$=\{(\pi(u),\pi(v)):(u,v)\in E\setminus E^{\prime}\}$
=$=\{(\pi(u),\pi(v)):(u,v)\in E\text{ and }(u,v)\notin E^{\prime}\}$
=$=\{(\pi(u),\pi(v)):(u,v)\in E\}\setminus\{(\pi(u),\pi(v)):(u,v)\in E^{\prime}\}$
@@
AddNode$(G,u)$ ${\mathrm{Orb}}(G,u)$
RemoveNode$(G,v)$ ${\mathrm{Orb}}(G,v)$
AddEdge$(G,u,v)$ ${\mathrm{Orb}}(G,u,v)$
RemoveEdge$(G,u,v)$ ${\mathrm{Orb}}(G,u,v)$
SetNodeAttribute$(G,u,t)$ ${\mathrm{Orb}}(G,u)$
SetEdgeAttribute$(G,u,v,t)$ ${\mathrm{Orb}}(G,u,v)$
SetGraphAttribute$(G,t)$ ${\mathrm{Orb}}(G,V)$
@@
Number of forward orbit-equivalent actionsNumber of backward orbit-equivalent $\frac{\text{Number of forward orbit-equivalent actions}}{\text{Number of backward orbit-equivalent actions}}=\frac{|{\mathrm{Orb}}(G,e)|}{|{\mathrm{Orb}}(G^{\prime},e)|}=\frac{|{\mathrm{Aut}}(G)|}{|{\mathrm{Aut}}(G^{\prime})|}.$
@@
For AddNodeAttribute$(G,u,t)$, the proof is straightforward: the only difference between $G$ and $G^{\prime}$ is the attribute assigned to node $u$, so it is immediate that ${\mathrm{Stab}}(G,u)={\mathrm{Stab}}(G^{\prime},u)$. The argument for AddEdgeAttribute is nearly identical. Next, we prove the corresponding result for the AddNode.
@@
=|¯$=|\bar{\mathcal{E}}(G)\cap a|\cdot p_{\bar{\mathcal{E}}}(e|G)$
for any $G\in s$. Since |¯$|\bar{\mathcal{E}}(G)\cap a|$ represents the number of orbit-equivalent actions from graph $G$, it is equal to $|{\mathrm{Orb}}(G,e)|$ for any $e\in\bar{\mathcal{E}}(G)\cap a$, where we defined the orbit of graph actions in Appendix F.4. Thus,

### v2

@@ 3.2 Generative Flow Networks
$\vskip 0.0pt{\mathcal{L}}_{\mathrm{TB}}(\tau)=\left(\log\frac{Z\prod_{t=0}^{n-1}p_{\mathcal{A}}(s_{t+1}|s_{t})}{R(s_{n})\prod_{t=0}^{n-1}q_{\mathcal{A}}(s_{t}|s_{t+1})}\right)^{2}.$
@@ 4.1 Problem Definition
Since isomorphism is an equivalence relation, it partitions the space $\mathcal{G}$} into classes, where each graph in a class is structurally identical to the others. Let $G]=\{G^{\prime}\in{\mathcal{G}}:G^{\prime}\cong G\}$ denote the equivalence class of $G$ induced by graph isomorphism. The state space $\mathcal{S}$} is defined as the set of equivalence classes of graphs, ${\mathcal{S}}=\{[G]:G\in{\mathcal{G}}\}$, rather than the graph space $\mathcal{G}$} itself. This is because our goal in using GFlowNets is to sample any graph within the equivalence class $s=[G]$ in proportion to $R(s)$. State transitions $\mathcal{A}$} can also be defined by partitioning the graph transitions $\mathcal{E}$} by the following equivalence relation:
@@
In practice, a graph generation process $({\mathcal{G}},{\mathcal{E}})$ is constructed by first designing a set of allowable actions in a given graph. For example, AddEdge(G,u,v)(G,u,v) adds an edge $(u,v)$ to the existing graph $G$, and $\texttt{AddNode}(G,u)$ adds a new node to an existing node $u$. The $\texttt{Stop}(G)$ terminates the process, in which case the graph-level attribute is flagged as terminated. Through this construction, graph transitions $\mathcal{E}$} are ensured to be structured so that any pair of isomorphic graphs have isomorphic successors.
@@
As an alternative, Ma et al. (2024) suggested using orbits to identify transition-equivalent actions, which we formalize next. Define graph actions as triples, $\bar{\mathcal{E}}=\{(G,t,u)\}$, where $t$ denotes the action type and $u$ specifies the nodes or edges affected by the action. Since graph transitions $\mathcal{E}$} are constructed by predefined set of graph actions, there is a one-to-one correspondence between $\mathcal{E}$} and $\bar{\mathcal{E}}$. However, we explicitly distinguish between them to introduce another equivalence relation.
@@
Given the distinction between transition-equivalent actions $\mathcal{A}$} and orbit-equivalent actions $\bar{\mathcal{A}}$, we define the state-action probability $p_{\bar{\mathcal{A}}}(a|s)$ differently from the state transition probability $p_{\mathcal{A}}(s^{\prime}|s)$: while $p_{\mathcal{A}}(s^{\prime}|s)$ accounts for all transition-equivalent actions, $p_{\bar{\mathcal{A}}}(a|s)$ aggregates only over orbit-equivalent actions. Formally, for a given graph $G$, let $\bar{\mathcal{E}}(G)$ denote the set of graph actions available from $G$. Then, for a state-action pair $s\in{\mathcal{S}}$ and $a\in\bar{\mathcal{A}}$, the state-action probability is defined as:
@@
where $p_{\bar{\mathcal{E}}}(e|G)$ denotes $p_{\mathcal{E}}(G^{\prime}|G)$ for $G^{\prime}$ being the next graph. Here, $\bar{\mathcal{E}}(G)\cap a$ represents the set of orbit-equivalent actions from $G$. By Theorem 4.3, it follows that $p_{\bar{\mathcal{A}}}(a|s)\leq p_{\mathcal{A}}(s^{\prime}|s)$ in general.
@@ 4.2 Properties of Equivalent Actions
Figure 2: Graphs representing transitions $(G_{1},G_{2},G_{3})$, where the first transition is performed by AddNode and the second by AddEdge. The number of forward/backward actions are represented as the number of arrows. Symmetries in each graph is related to orbit-equivalent actions, as seen in the ratio $|{\mathrm{Aut}}(G_{1})|/|{\mathrm{Aut}}(G_{2})|=|{\mathrm{Orb}}(G_{1},$ $)|/|{\mathrm{Orb}}(G_{2},$ )|)|. Nodes in the same orbit are given the same color.
@@
where $|\bar{\mathcal{E}}(G)\cap a|$ represents the number of orbit-equivalent actions. This is because when actions are parameterized using graph neural networks (GNNs), orbit-equivalent actions are assigned equal probabilities. In general, permutation-equivariant functions produce identical representations for nodes within the same orbit (as detailed in Appendix E.1). When node representations are aggregated to compute edge representations using invariant aggregators such as SUM or MEAN, edges within the same orbit also receive identical representations. Alternative parameterizations, such as the relative edge parameterization proposed by Shen et al. (2023), also assign equal probabilities to orbit-equivalent actions, while enhancing representational power.
Note that the number of orbit-equivalent actions, $|\bar{\mathcal{E}}(G)\cap a|$, corresponds to the size of the orbit of the associated nodes or edges. The following lemma shows that, when considering ratios, counting orbit-equivalent actions simplifies to counting automorphisms.
@@
We presented Lemma 4.5 in the context of AddEdge for simplicity, but similar lemma holds for other action types such as AddNode, AddNodeAttributes, and others used in node-by-node generation (See Appendix C for the full list of considered actions). We also consider AddFragment in our experiments, but discuss its properties separately in Appendix H, as the corresponding formula differs in fragment-based schemes.
@@ 5 Symmetry-Aware GFlowNets
Assume that $G_{0}$ is the empty graph or a single node, so that $|{\mathrm{Aut}}(G_{0})|=1$. Given the complete graph trajectory $\tau=(G_{0},G_{1},\dots,G_{n})$, constructed using a node-by-node generation scheme, the trajectory balance loss is given by:
@@
Consider a node-by-node graph generation scheme. We define the graph-level detailed balance condition, as opposed to the standard state-level condition, as follows:
@@
Together with Corollary 5.1, we see that scaling the reward alone is sufficient for both TB and DB objectives. This suggests that other GFlowNet objectives, such as subtrajectory balance (Madan et al., 2023) and flow-matching (Bengio et al., 2021), can also be used with reward scaling (see the discussion on the flow-matching Appendix G). This provides a straightforward approach to implementing GFlowNet objectives while reducing the computational burden of counting automorphisms at each transition.
@@
Let $G$ represents a terminal state ($G]\in{\mathcal{X}}$) generated by connecting $k$ fragments $\{C_{1},\dots,C_{k}\}$. Then, the scaled rewards to offset the effects of equivalent actions are given by:
@@
Relation to distribution learning
While our work focuses on improving the reward-matching capabilities of GFlowNets, many graph generation methods instead aim to learn the data distribution by maximizing a variational lower bound (VLB):
$\begin{split}\log\bar{p}_{{\mathcal{A}}}(x)\geq\mathbb{E}_{\tau\sim q(\tau|x)}\left[\log p_{\mathcal{A}}(\tau)-\log q(\tau|x)\right].\end{split}$ (5)
Here, $q(\tau|x)$ is a variational distribution that samples trajectories $\tau=(s_{0},\dots,s_{n}=x)$ conditioned on the final state $x$. There are two primary ways to parameterize $q$ and sample trajectories $\tau$: (1) $q$ can be implemented as a backward policy $q(s_{t}|s_{t+1})$, as in GFlowNets; (2) $q(\pi|x)$ can first sample a node or edge ordering $\pi$, which then deterministically defines a trajectory $\tau$. Our method suggest that when the backward policy is used to define $q(s_{t}|s_{t+1})$, the VLB can be computed via automorphism counting. In contrast, when $q(\pi|x)$ is used to sample orderings, no further correction is required. This is because forward equivalent actions split probabilities across different orderings that induce the same trajectory. Since this result differs from that of Chen et al. (2021), we elaborate on this point in Appendix I, where we interpret equivalent actions as those that correspond to different orderings but ultimately lead to the same state sequence.
@@ 6 Experiments
In this section, we conduct experiments to validate our theoretical results and demonstrate the effectiveness of our method. We use a uniform backward policy across all experiments. Details on hyperparameters and model configurations can be found in Appendix K. The experiments compare the following methods: (1) Vanilla GFlowNets, which do not incorporate graph symmetries. (2) Transition Correction, which identifies transition-equivalent actions by performing multiple isomorphism tests and sums their probabilities accordingly. (3) PE, the method proposed by Ma et al. (2024), which approximately identifies orbit-equivalent actions using positional encoding. (4) Reward Scaling, which achieves correction by modifying only the reward signals. (5) Flow Scaling, which multiplies symmetry ratio $|{\mathrm{Aut}}(G^{\prime})|/|{\mathrm{Aut}}(G)|$ to backward probability at each transition. Note that methods 3-5 are orbit correction methods. Since Reward Scaling and Flow Scaling have the same effect under the TB objective, we only consider Flow Scaling when using the DB objective. [^1] Source code available at: https://github.com/hohyun312/sagfn
@@ 6.3 Molecule Generation
[…]
@@ 7 Discussion and Conclusion
In this paper, we analyzed the properties of equivalent actions and proposed a simple correction method that allows for unbiased sampling from the target distribution. Our analysis shows that, without correction, highly symmetric graphs are less likely to be sampled, while symmetric fragments are more likely to be sampled, which is crucial for molecule discovery. We demonstrated that the reward-scaling technique works for both TB and DB objectives. Experimental results suggest that reward scaling and flow scaling effectively removes bias, allowing for accurate modeling of the target distribution, which is essential for sampling high-reward molecules. The exact effect, however, depends on the reward structure of the given task.
@@ Appendix C Definitions of Graph Actions
  • AddNode(G,u)(G,u) adds a new node to the existing node $u$.
  • AddEdge(G,u,v)(G,u,v) adds a new edge $(u,v)$.
  • AddFragment(G,C)(G,C) adds a fragment $C$.
  • RemoveNode(G,v)(G,v) removes node $v$ and its connecting edges.
  • RemoveEdge(G,u,v)(G,u,v) removes the edge $(u,v)$.
  • RemoveFragment(G,C)(G,C) removes the subgraph $C$.
  • SetNodeAttribute(G,u,t)(G,u,t) sets the node-level attribute $t$ for node $u$.
  • SetEdgeAttribute(G,u,v,t)(G,u,v,t) sets the edge-level attribute $t$ for the edge $(u,v)$.
  • SetGraphAttribute(G,t)(G,t) sets a graph-level attribute $t$.
@@ F.1 Proof of Equation 6
$\pi(E\cup E^{\prime})$ $=\{(\pi(u),\pi(v)):(u,v)\in E\cup E^{\prime}\}$
$=\{(\pi(u),\pi(v)):(u,v)\in E\text{ or }(u,v)\in E^{\prime}\}$
$=\{(\pi(u),\pi(v)):(u,v)\in E\}\cup\{(\pi(u),\pi(v)):(u,v)\in E^{\prime}\}$
@@
$\pi(E\setminus E^{\prime})$ $=\{(\pi(u),\pi(v)):(u,v)\in E\setminus E^{\prime}\}$
$=\{(\pi(u),\pi(v)):(u,v)\in E\text{ and }(u,v)\notin E^{\prime}\}$
$=\{(\pi(u),\pi(v)):(u,v)\in E\}\setminus\{(\pi(u),\pi(v)):(u,v)\in E^{\prime}\}$
@@
AddNode(G,u)(G,u) ${\mathrm{Orb}}(G,u)$
RemoveNode(G,v)(G,v) ${\mathrm{Orb}}(G,v)$
AddEdge(G,u,v)(G,u,v) ${\mathrm{Orb}}(G,u,v)$
RemoveEdge(G,u,v)(G,u,v) ${\mathrm{Orb}}(G,u,v)$
SetNodeAttribute(G,u,t)(G,u,t) ${\mathrm{Orb}}(G,u)$
SetEdgeAttribute(G,u,v,t)(G,u,v,t) ${\mathrm{Orb}}(G,u,v)$
SetGraphAttribute(G,t)(G,t) ${\mathrm{Orb}}(G,V)$
@@
[…]
@@
For AddNodeAttribute(G,u,t)(G,u,t), the proof is straightforward: the only difference between $G$ and $G^{\prime}$ is the attribute assigned to node $u$, so it is immediate that ${\mathrm{Stab}}(G,u)={\mathrm{Stab}}(G^{\prime},u)$. The argument for AddEdgeAttribute is nearly identical. Next, we prove the corresponding result for the AddNode.
@@
$=|\bar{\mathcal{E}}(G)\cap a|\cdot p_{\bar{\mathcal{E}}}(e|G)$
for any $G\in s$. Since $|\bar{\mathcal{E}}(G)\cap a|$ represents the number of orbit-equivalent actions from graph $G$, it is equal to $|{\mathrm{Orb}}(G,e)|$ for any $e\in\bar{\mathcal{E}}(G)\cap a$, where we defined the orbit of graph actions in Appendix F.4. Thus,